OpenStudy (anonymous):

Help please!! Trigonometric substitution integrate 1/(x^5*sqrt((9x^2-1)) on interval [sqrt(2)/3,2/3]

OpenStudy (anonymous):

You wanna use \(x=\sec(\theta)/3\)

OpenStudy (anonymous):

thanks for responding.. why is it 1/3 and not just 3?

OpenStudy (anonymous):

Remember:\[ b^2x+a^2 \to x = \frac{a}{b}\tan\theta \\ b^2x-a^2 \to x = \frac{a}{b}\sec\theta \\ a^2-b^2x \to x = \frac{a}{b}\sin\theta \]

OpenStudy (anonymous):

In this case \(b^2=9\) and \(a^2=1\), this means \(a/b = 1/3\)

OpenStudy (anonymous):

You want to have \((1/3)^2\) to cancel out the \(9\) coefficient next to \(x\).

OpenStudy (anonymous):

Ohhh, that I didn't know. Thank you.

OpenStudy (anonymous):

and then my boundaries would change to pi/18 and ??

OpenStudy (anonymous):

Well remember that \[ x = \frac{\sec(\theta)}{3} = \frac{1}{3\cos(\theta)}\implies \frac{1}{\cos(\theta)} = 3x\implies \cos\theta = \frac{1}{3x} \]

OpenStudy (anonymous):

So you have \(\cos(\theta_1) =1/\sqrt{2} \), \(\cos(\theta_2) =1/2 \)

OpenStudy (anonymous):

These are fairly common angles.

OpenStudy (anonymous):

ok, i'll give it another shot. thanks for your help

OpenStudy (anonymous):

Got it! Really appreciate your help!